In a recent year, some professional baseball players complained that umpires were calling more strikes than the average rate of 61.0% called the previous year. At one point in the season, umpire Dan Morrison called strikes in 2231 of 3581 pitches.
This is a hypothesis question. I am really struggling on how to solve these types of questions, if someone can show me step by step how it works with the answer I can do the other 32 questions I have, I just can't seem to get it started.

Answer by stanbon(75887)

In a recent year, some professional baseball players complained that umpires were calling more strikes than the average rate of 61.0% called the previous year. At one point in the season, umpire Dan Morrison called strikes in 2231 of 3581 pitches.
--------------------
Note: There are hundreds of examples of worked statistics problems on
this site.
--------------------
Ho: p = 0.61
Ha: p > 0.61
---
sample proportion = phat = 2231/3581 = 0.623
-------------------
Critical Value for right-tail test with alpha = 5%: 1.645
Test Statistic: z(0.623) = (0.623-0.61)/sqrt(0.61*0.39/3581) = 1.5962
p-value = P(z> 1.5962) = 0.0552
-------------------
Conclusion: Since the p-value is greater than 5%, fail to reject Ho.
Probably you should do more testing also as the decision evidence is
weak.
-------------------
Cheers,
Stan H.